OpenStudy (anonymous):

Give the acceleration a = d^2s/dt^2, initial velocity, and initial position of a body moving on a coordinate line. Find the body's position at time t. a = -4 sin 2t, v(0)=2, s(0)=-3

OpenStudy (anonymous):

\[a(t)=s''(t)\]\[v(t)=s'(t)\]using above, the equation can be written\[s''=-4 \sin 2t\]\[s'(0)=2\]\[s(0)=-3\]solve the second order DE

OpenStudy (dumbcow):

in other words...integrate function twice, using the initial values to solve for constants of integration \[\int\limits -4\sin(2t) dt = 2\cos(2t) +C\] --> 2cos(0) +C = 2 --> C = 0 \[\int\limits 2 \cos(2t)dt = \sin(2t) +C\] --> sin(0) +C = -3 ----> C = -3 \[s(t) = \sin(2t) -3\]